OpenStudy (itskayylayy1213):

8(2x – 14) + 13 = 4x – 27 Part B: What value of x makes the equation true? (4 points) i got part a but i dont understand part b. the answer is 6 if that helps.

Mehek (mehek14):

8(2x – 14) + 13 = 4x – 27 16x - 112 + 13 = 4x - 27 16x -99 = 4x - 27 12x = 72 72/12 = ?

OpenStudy (itskayylayy1213):

8 (2x - 14) + 13 = 4x -27 distribute 8 through the parentheses 16x - 112 + 13 = 4x - 27 combine like terms 16x - 99 = 4x -27 subtract 4x from each side 12x - 99 = -27 add 99 to each side 12x = 72 divide each side by 12 x = 6

Mehek (mehek14):

Correct, so the value of x that makes it true is 6

OpenStudy (lebribacktho_16):

yeah its 6 Lmao

OpenStudy (itskayylayy1213):

oh that makes sense x'd thanks

OpenStudy (itskayylayy1213):

ik its 6

Mehek (mehek14):

Then what do you need help with?

Mehek (mehek14):

part b is x = 6

OpenStudy (itskayylayy1213):

Part B: What value of x makes the equation true? thats what i needed help on bc my brain is slow but i got it now

Mehek (mehek14):

. . . If we solve the equation, we get x = 6 that means the value of x that makes the equation true is 6

OpenStudy (flashp):

Part B) x = 6 makes the equation true